OpenStudy (anonymous):

Find the quotient and express the answer in scientific notation. 302 (9.1 x 10^4) A) 3.32 x 10-4 B) 3.32 x 10-3 C) 3.32 x 104 D) 3.32 x 103

OpenStudy (anonymous):

\[302*9.1*10^4=2748.2*10^4=2.7482*10^7 \]